Shane's neighbor pledged $1.25 for every 0.5 miles that Shane swims in the charity swim-a-thon. If Shane swims 3 miles, how much money will his neighbor donate?
Answer by stanbon(75887)

Shane's neighbor pledged $1.25 for every 0.5 miles that Shane swims in the charity swim-a-thon. If Shane swims 3 miles, how much money will his neighbor donate?
---
3 miles = 6*0.5 miles
---
Donation = 6*1.25 = $7.50
=============================
Cheers,
Stan H.
=============